# Data-Driven Logistic Regression Ensembles With Applications in Genomics

### v5

@@ Abstract
Advances in data collecting technologies in genomics have significantly increased the need for tools designed to study the genetic basis of many diseases. Statistical tools used to discover patterns between the expression of certain genes and the presence of diseases should ideally perform well in terms of both prediction accuracy and identification of key biomarkers. We propose a new approach for dealing with high-dimensional binary classification problems that combines ideas from regularization and ensembling. The ensembles are comprised of a relatively small number of highly accurate and interpretable models that are learned directly from the data by minimizing a global objective function. We derive the asymptotic properties of our method and develop an efficient algorithm to compute the ensembles. We demonstrate the good performance of our method in terms of prediction accuracy and identification of key biomarkers using several medical genomics datasets involving common diseases such as cancer, multiple sclerosis and psoriasis. In several applications our method could identify key biomarkers that were absent in state-of-the-art competitor methods. We develop a variable importance ranking tool that may guide the focus of researchers on the most promising genes. Based on numerical experiments we provide guidelines for the choice of the number of models in our ensembles.
@@ 1 Introduction
We propose a new approach to learn a diverse ensemble of sparse classification models that is especially well suited for high-dimensional medical genomics applications. We demonstrate the advantage of our ensemble method in terms of both prediction and identification of key biomarkers using several datasets from the GEO database. These datasets involve different types of common diseases such as cancer, multiple sclerosis and psoriasis. In particular, we show that the prediction accuracy of the sparse models in our ensembles matches the prediction accuracy of standard single-model sparse methods. Since the models in the ensembles are learned simultaneously and directly form the data (free of randomization or other heuristics) by optimizing a global objective function, they each provide an alternative explanation for the relationship between the genes and disease. The objective function contains a penalty that promotes diversity in the obtained models so that individual models may be driven by potentially different biological mechanisms. With respect to the identification of important predictor genes, our examples demonstrate that the proposed ensemble method can identify key biomarkers that are discarded by state-of-the-art sparse methods and ensemble variable ranking methods. At the same time, we show that our method generally includes key biomarkers selected by sparse methods or flagged as important by ensemble variable ranking methods.
@@
The remainder of the paper is organized as follows. In Section 2 we briefly review the literature on sparse and ensemble methods, and we analyze lung and thyroid cancer genomic datasets to highlight the potential benefit of our hybrid approach over standard state-of-the-art sparse and ensemble methods. In Section 3 we introduce our method to learn a data-driven diverse ensemble of sparse logistic regression models. We show that this new ensemble method is consistent under mild regularity assumptions. We illustrate how the proposed method can uncover multiple mechanisms for the relationship between predictors and the response on the thyroid cancer genomics dataset. An efficient block coordinate descent algorithm to solve the multi-convex optimization problem is provided in Section 4. Section 5 presents the results of an extensive simulation study to compare our method with a large number of state-of-the-art competitors for high-dimensional binary classification. In Section 6 we further show the excellent performance of the proposed method in a numerical study using a large collection of high-dimensional medical genomic datasets. We also show how our method can be used to rank genes in order of importance. To provide insight in how to choose the number of models in our method, we study the accuracy-diversity trade-off and computational cost of our method in Section 7. Potential directions for further research are provided in Section 8.
2 Empirical Motivation
In a typical medical genomics data application the training data consist of thousands of gene expression level measurements on a limited number of samples, collected from some cancer tumor tissue or from some adjacent normal tissue for example. In such an application the main goal is to train a classification model that accurately identifies the tumor status of the tissue samples using the available gene expression data. At the same time, medical researchers wish to identify important subsets of differentially expressed genes and the way they operate.
As motivating examples we consider the lung and thyroid cancer genomic datasets, GSE10245 and GSE5364 respectively, from the Gene Expression Omnibus (GEO) database. GSE10245 consists of $n=58$ non-small cell lung cancer (NSCLC) tissue samples of subtypes adenocarcinoma (AC) or squamous cell carcinoma (SCC) with $p=54,675$ gene expression measurements. GSE5364 consists of $n=51$ cell tissue samples of either thyroid cancer tumors or adjacent normal tissues with $p=22,283$ gene expression measurements. The objective is to construct a classifier that can accurately specify the cancer tumor status of the samples, AC or SCC subtype for the lung cancer dataset and cancer tumor or adjacent tissue sample for the thyroid cancer dataset, using their gene expression levels. The classifier should also identify, for each dataset, all possible relevant biomarkers that play a role in the differentiation between the two types of cell tissue samples.
The data are pre-processed in a standard way by using a three step procedure similar to Dudoit et al. (2002): (1) thresholding of gene expression levels with a floor of 20 and a ceiling of 20,000; (2) exclusion of genes for which the ratio between the maximum and minimum expression level is less than 2; (3) exclusion of genes for which the difference between the maximum and minimum expression level is less than 200; (4) base-2 logarithmic transformation. For most methods we then retain from the remaining genes the $p=500$ genes which are most important according to their $q$-value (Storey, 2002; Storey et al., 2023). The $q$-value is a popular alternative to the $p$-value for high-dimensional data that controls for the false discovery rate and alleviates the issue of overestimating significance that arises from using multiple pairwise $t$-tests.
Training a classification model for these examples is particularly challenging due to the large number of gene expression measurements relative to the available number of tissue samples. The two main approaches proposed in the literature for dealing with such high-dimensional data are sparse methods and ensemble methods, respectively. We briefly review some state-of-the-art methods from both approaches and highlight their strengths and limitations on these datasets. We also use these examples to illustrate the good performance of our hybrid method, which uses an ensemble of sparse models as explained in Section 3, both in terms of prediction accuracy and identification of target biomarkers.
@@ 2.1 Sparse Methods
Let $\mathcal{D}=\left(\mathbf{y},\mathbf{X}\right)$ denote the training data, where $\mathbf{y}\in\mathbb{R}^{n}$ is the vector of class labels and $\mathbf{X}\in\mathbb{R}^{n\times p}$ is the design matrix which consists of $n$ measurements $\mathbf{x}_{i}$ on $p$ predictors. The classes are labeled by $y_{i}\in\{-1,1\}$ for convenience and the predictor variables have been standardized, i.e. $\sum\limits_{i=1}^{n}x_{ij}/n=0$ and $\sum\limits_{i=1}^{n}x_{ij}^{2}/n=1$, $1\leq j\leq p$. Logistic regression models the class-conditional probabilities through a nonlinear function of the predictor variables,
@@
where $\beta_{0}$ and $\boldsymbol{\beta}\in\mathbb{R}^{p}$ are the intercept and vector of regression coefficients, respectively. The function $S(t)=e^{t}/(1+e^{t})$ is the well-known logistic function. With $f(\mathbf{x}_{i})=\beta_{0}+\mathbf{x}_{i}^{T}\boldsymbol{\beta}$, let
$\mathcal{L}\left(f(\mathbf{x}_{i}),y_{i}\right)=\mathcal{L}(\beta_{0},\boldsymbol{\beta}\mid y_{i},\mathbf{x}_{i})=\log\left(1+e^{-y_{i}f(\mathbf{x}_{i})}\right),$ (2)
denote the logistic regression loss function. Minimizing the corresponding empirical loss
$\mathcal{V}_{n}(f)=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}\left(f(\mathbf{x}_{i}),y_{i}\right),$ (3)
leads to overfitting of the training data and poor out-of-sample prediction accuracy when $p\gg n$.
@@
$\min_{\beta_{0}\in\mathbb{R},\,\boldsymbol{\beta}\in\mathbb{R}^{p}}\mathcal{V}_{n}(f)+\lambda P\left(\boldsymbol{\beta}\right)$ (4)
[…]
@@ 2.2 Ensemble Methods
The need for diversity among the models in an ensemble has been studied more broadly, with random forests (Breiman, 2001) being a very popular example. In random forests the individual decision trees are usually weakened by random selection of candidate features at each node, resulting in de-correlated trees that strengthen the ensemble accuracy. Similarly, in the random generalized linear models (RGLM) method of Song et al. (2013), random sampling of the data (bagging) (Breiman, 1996) and the random predictor subspace method (Ho, 1998) are combined to create varying samples and predictor sets, resulting in weak but de-correlated individual models. In the (extreme) gradient boosting method of Chen and Guestrin (2016), diverse trees are generated by sequentially fitting the residuals of the previous fit rather than the original responses, resulting in a collection of trees that are poor individual predictors but can produce ensembles with high accuracy.
In summary, current ensemble methods lack interpretability as they generally generate many weak models that are only useful if a large number of them are pooled together. However, ad hoc tools have been introduced in the literature to assess predictor importance in some blackbox ensemble methods such as the Gini importance index in random forests (Breiman, 2001). Finally, due to the heuristic nature of current ensemble methodology, less theory has been developed to study their asymptotic behavior.
2.3 Methods and Performance Measures
We use the thyroid (GSE5364) and lung cancer (GSE10245) genomic datasets to illustrate the performance of several existing state-of-the-art sparse and ensemble methods and to compare them to our new proposal. We use all methods as implemented in the CRAN (R Core Team, 2022) R packages given below:
  1. 1–2.
    Lasso and Elastic Net (EN) for logistic regression from the glmnet package (Friedman et al., 2010).
  2. 3–4.
    Adaptive Lasso and Relaxed Lasso for logistic regression from the gcdnet (Yang and Zou, 2017) and glmnet packages, respectively.
  3. 5.
    Minimum concave penalized (MCP) logistic regression from the ncvreg package (Breheny and Huang, 2011).
  4. 6.
    Sure Independence Screening (SIS) with the SCAD penalty from the SIS package (Saldana and Feng, 2018).
  5. 7–8.
    Split-Lasso and Split-EN logistic regression which will be introduced in the following sections as implemented in the SplitGLM package (Christidis et al., 2021).
  6. 9.
    Random GLM (RGLM) from the RGLM package (Song and Langfelder, 2013).
  7. 10.
    Random Forest (RF) from the randomForest package (Liaw and Wiener, 2002).
  8. 11.
    Extreme Gradient Boosting (XGB) from the xgboost package (Chen et al., 2020).
[…]
To evaluate prediction performance we report the misclassification rate (MR) of all the methods. For the ensemble methods (Split-Lasso, Split-EN, RGLM, RF and XGB) we also report the average misclassification rate of the individual models ($\overline{\text{MR}}$) which measures the prediction accuracy of the individual models used in an ensemble. To estimate MR and $\overline{\text{MR}}$, we randomly split the lung and thyroid datasets $N=50$ times into training sets of size $n=29$ and $n=26$, respectively, to fit the models, and corresponding tests sets of size $m=29$ and $m=25$, respectively, to evaluate the prediction performance of the methods.
To evaluate the performance of the methods with respect to identification of target biomarkers, we investigate the inclusion or absence of several key genes identified in the literature for lung and thyroid cancer in the resulting models.
2.4 Prediction Performance
In Table 1 we report the MR of the sparse and ensemble methods averaged over the $N=50$ random splits of the lung and thyroid cancer genomic datasets. For the ensemble methods we also report the average individual model $\overline{\text{MR}}$. In each column the best performances are marked in bold. If we leave the split logistic regression methods out of the comparison, then RGLM with $G=100$ and $G=200$ models yields the best performance in terms of MR for both datasets. For the thyroid cancer dataset, also RF with $G=500$ and $1,000$ models yields the best performances. However, the average individual model $\overline{\text{MR}}$ reveals that at the same time the individual models in RF and RGLM yield a weak performance. Moreover, while RF and RGLM with a large number models perform very well, their performance deteriorates when a small number of models is used. In RF and RGLM diversity between the individual models is achieved using randomization which results in uninterpretable, weak individual models. This approach requires a large number of decorrelated models to achieve a low MR.
Table 1: Average MR and $\overline{\text{MR}}$ of sparse and ensemble methods estimated using $N=50$ random splits of the lung (GSE10245) and thyroid (GSE5364) cancer genomic datasets into training sets with $n=29$ and $n=26$ samples respectively, and test sets with the remaining samples. The standard errors are in parenthesis.
[…]
=2pt GSE10245 GSE5364 Method MR $\mathbf{\overline{MR}}$ MR $\mathbf{\overline{MR}}$ Lasso 0.09 (0.05) -- 0.13 (0.07) -- EN 0.08 (0.05) -- 0.11 (0.06) -- Adaptive 0.15 (0.08) -- 0.14 (0.10) -- Relaxed 0.09 (0.05) -- 0.14 (0.08) -- MCP 0.16 (0.09) -- 0.18 (0.11) -- SIS-SCAD 0.16 (0.08) -- 0.19 (0.11) -- Split-Lasso-5 0.08 (0.05) 0.09 (0.04) 0.10 (0.06) 0.14 (0.06) Split-Lasso-10 0.07 (0.04) 0.09 (0.03) 0.10 (0.05) 0.15 (0.05) Split-Lasso-25 0.07 (0.05) 0.10 (0.03) 0.09 (0.05) 0.16 (0.04) Split-EN-5 0.07 (0.04) 0.08 (0.04) 0.10 (0.05) 0.12 (0.05) Split-EN-10 0.07 (0.04) 0.08 (0.03) 0.10 (0.05) 0.13 (0.05) Split-EN-25 0.07 (0.04) 0.09 (0.03) 0.09 (0.05) 0.14 (0.04) RGLM-5 0.10 (0.05) 0.48 (0.08) 0.15 (0.09) 0.55 (0.09) RGLM-10 0.09 (0.05) 0.49 (0.07) 0.12 (0.07) 0.54 (0.07) RGLM-25 0.09 (0.05) 0.49 (0.05) 0.11 (0.07) 0.55 (0.05) RGLM-100 0.08 (0.05) 0.49 (0.04) 0.10 (0.07) 0.55 (0.04) RGLM-200 0.08 (0.05) 0.48 (0.04) 0.09 (0.07) 0.55 (0.04) RF-5 0.15 (0.07) 0.25 (0.06) 0.21 (0.08) 0.30 (0.07) RF-10 0.12 (0.06) 0.25 (0.04) 0.16 (0.08) 0.30 (0.05) RF-25 0.11 (0.07) 0.25 (0.03) 0.11 (0.07) 0.30 (0.03) RF-500 0.10 (0.07) 0.25 (0.02) 0.10 (0.08) 0.31 (0.02) RF-1,000 0.10 (0.08) 0.25 (0.02) 0.09 (0.07) 0.30 (0.02) XGB 0.19 (0.09) 0.25 (0.07) 0.19 (0.12) 0.22 (0.10)
As will be seen in Section 3, split logistic regression ensembles are modeled directly from the data without relying on randomization or any heuristics. Moreover, the degree to which the models are sparse and diverse is driven directly by the data. For both the lung and thyroid cancer genomic datasets, the split logistic regression methods obtained sparse and diverse individual models whose prediction accuracy $\overline{\text{MR}}$ either match or nearly match the performance of the corresponding single model sparse estimator. Hence, each of the individual models in Split-Lasso/Split-EN is just as reliable as the single-model sparse Lasso/EN estimators. Hence, each individual model in split logistic regression provides a potential explanation for the relationship between the genetic covariates and the cancer diagnosis of the patient. By ensembling a relatively small number of these accurate, sparse and diverse models, the performance of split logistic regression ensembles is comparable to the performance of the best blackbox ensemble methods. The overall good predictive performance of split logistic regression compared to state-of-the-art competitors is shown further in Section 6 by a benchmark study using a large collection of high-dimensional genomics datasets.
Beyond the good predictive performance of split logistic regression ensembles and the individual models that comprise them, they can be used to uncover potentially different mechanisms that explain the relationship between the gene expression levels and the tumor status of the tissue samples. Over the $N=50$ random splits, the individual models in Split-EN ensembles with $G=10$ were comprised of 23 genes on average for both the lung and thyroid cancer datasets. In comparison, the logistic elastic net contained on average 44 and 38 genes, respectively. Since split logistic regression learns the ensembles directly from the data, the inclusion of a gene in any of the models is the result of its contribution to the minimization of the global objective function which includes a sparsity and a diversity penalty. Thus, each model in the ensemble is a sparse, fully interpretable model that can reveal the relation between the tumor status and a set of genes.
2.5 Identification of Target Biomarkers
An advantage of split logistic regression ensembles over their single model counterparts is that by combining information from several individual models, they may identify more relevant genes. For example, over the $N=50$ random splits of the lung cancer genomic dataset, the Split-EN ensembles based on $G=10$ models contained on average 231 genes compared to only 44 genes for the logistic elastic net. Since the inclusion of a gene in the Split-EN ensemble is due to its importance in the minimization of its objective function, it may be expected that these genes are relevant to differentiate between the AC and SCC lung cancer subtype. For example, both Split-EN with $G=10$ models and the logistic elastic net included the gene CGN in all $N=50$ random splits of the lung cancer dataset, a biomarker that has been identified as highly expressed in AC but significantly reduced in SCC (Paschoud et al., 2007). Also the genes PTGFRN, SOX2 and TMC5 have all been identified as key biomarkers to differentiate between AC and SCC cancer subtypes (see George et al., 2015; Karachaliou et al., 2013; Xiao et al., 2017, respectively). Split-EN with $G=10$ models indeed includes these genes in the model more than 80% of the times, while they were included in less than 30% of the logistic elastic net models over the same random splits of the data. We also note that the genes selected by the logistic elastic net were always a subset of the genes selected by the Split-EN ensemble model with $G=10$ models. Similarly, genes that were consistently ranked as one of the top 250 most important genes by RF based on the Gini index were often also selected by Split-EN with $G=10$ models. For example, the gene IRF6 which has been found to drive the progression of NSCLC (Liu et al., 2021) was 88% of the times in the top 250 for RF. This gene was also included in 74% of Split-EN models with $G=10$. On the other hand, several of the genes most often selected by Split-EN such as the key biomarkers CGN, PTGFRN, SOX2 and TMC5 mentioned above, were never included in the top 250 by RF.
In split logistic regression genes may appear in more than one individual model, albeit at a cost in the global objective function (see Section 3). Hence, genes appearing in multiple individual models are expected to be more important biomarkers than others in terms of predicting the response. This information can thus be exploited to rank the genes in order of importance as explained in Section 6 based on the thyroid cancer dataset.
3 Split Logistic Regression Models

### v6

@@ Abstract
Advances in data collecting technologies in genomics have significantly increased the need for tools designed to study the genetic basis of many diseases. Effective statistical methods should excel in both prediction accuracy and biomarker identification. We introduce a novel approach to high-dimensional binary classification that integrates regularization with ensembling techniques. Our method constructs compact ensembles of interpretable models derived by optimizing a global objective function. In medical genomics applications, our approach identifies critical biomarkers overlooked by competing methods. We develop a variable importance ranking system to help researchers prioritize promising genes. The method’s asymptotic properties are established, and an efficient computational algorithm is provided. Through extensive simulations across complex scenarios and analysis of genomics datasets for cancer, multiple sclerosis, and psoriasis, we demonstrate strong predictive performance. Based on our numerical experiments, we offer practical guidelines for determining optimal ensemble size.
@@ 1 Introduction
We propose a new approach to learn a diverse ensemble of sparse classification models that is especially well suited for high-dimensional medical genomics applications. Specifically, we extend the regression ensemble method proposed by Christidis et al. (2020) to the binary classification setting. We demonstrate the advantage of our ensemble method in terms of both prediction and identification of key biomarkers using several datasets from the GEO database. These datasets involve different types of common diseases such as cancer, multiple sclerosis and psoriasis. In particular, we show that the prediction accuracy of the sparse models in our ensembles matches the prediction accuracy of standard single-model sparse methods. Since the models in the ensembles are learned simultaneously and directly form the data (free of randomization or other heuristics) by optimizing a global objective function, they each provide an alternative explanation for the relationship between the genes and disease. The objective function contains a penalty that promotes diversity in the obtained models so that individual models may be driven by potentially different biological mechanisms. With respect to the identification of important predictor genes, our examples demonstrate that the proposed ensemble method can identify key biomarkers that are discarded by state-of-the-art sparse methods and ensemble variable ranking methods. At the same time, we show that the proposed method generally includes key biomarkers selected by sparse methods or flagged as important by ensemble variable ranking methods.
@@
Following the methodological development framework of Heinze et al. (2024), this work encompasses phases I and II, introducing a novel method with initial validation through comprehensive simulations across various scenarios (sample sizes, sparsity, correlation structures, class imbalance, non-linearity, and interactions).While the results of our numerical experiments and medical genomics data applications demonstrate the promising performance of the proposed classification ensembles across a variety of data scenarios, we present this work as a method development study with strong proof-of-concept validation, rather than a claim of universal superiority. The success observed across our diverse testing scenarios suggests that phases III and IV—independent validation and practical implementation—may naturally follow as the method gains adoption in the research community.
@@
The remainder of the paper is organized as follows. In Section 2 we briefly review the literature on sparse and ensemble methods, and introduce the proposed method to learn a data-driven diverse ensemble of sparse logistic regression models. We show that this new ensemble method is consistent under mild regularity assumptions. In Section 3, we analyze lung and thyroid cancer genomic datasets to highlight the potential benefits of our hybrid approach over standard sparse and ensemble methods. We show how the method can uncover multiple mechanisms linking predictors to the response in the thyroid cancer dataset and how it can rank genes by importance. In Section 4 an efficient block coordinate descent algorithm to solve the multi-convex optimization problem is provided. Section 5 presents the results of an extensive simulation study to compare the proposed method with a large number of state-of-the-art competitors for high-dimensional binary classification. In Section 6 we further show the excellent performance of the proposed method in a numerical study using a large collection of high-dimensional medical genomic datasets. To provide insight in how to choose the number of models in the proposed method, we study the accuracy-diversity trade-off and computational cost of our method in Section 7. Potential directions for further research are provided in Section 8.
2 Split Logistic Regression
Training classification models on high-dimensional data, such as gene expression profiles where the number of measurements ($p$) exceeds the number of tissue samples ($n$), presents significant statistical and computational challenges. This $p>n$ scenario is common in bioinformatics and other fields. To address these challenges, two primary families of methods have emerged as dominant strategies in the literature: sparse modeling, which aims for feature selection and interpretability, and ensemble methods, which typically focus on maximizing predictive accuracy and stability by combining multiple models.
In this section, we briefly review representative state-of-the-art techniques from both families, discussing their strengths and limitations within the high-dimensional setting. Building upon this context, we then introduce Split Logistic Regression, our proposed hybrid approach specifically designed to combine the variable selection benefits of sparse methods with the generally superior predictive performance of ensemble learning.
@@ 2.1 Sparse Methods
Let $\mathcal{D}=\left(\mathbf{y},\mathbf{X}\right)$ denote the training data, where $\mathbf{y}\in\mathbb{R}^{n}$ is the vector of class labels and $\mathbf{X}\in\mathbb{R}^{n\times p}$ is the design matrix which consists of $n$ measurements $\mathbf{x}_{i}$ on $p$ predictors. We consider the binary classification problem where the classes are labeled by $y_{i}\in\{0,1\}$. The predictor variables have been standardized, i.e. $\sum\limits_{i=1}^{n}x_{ij}/n=0$ and $\sum\limits_{i=1}^{n}x_{ij}^{2}/n=1$, $1\leq j\leq p$. Logistic regression models the class-conditional probabilities through a non-linear function of the predictor variables,
@@
where $\beta_{0}$ and $\boldsymbol{\beta}\in\mathbb{R}^{p}$ are the intercept and vector of regression coefficients, respectively. The function $S(t)=e^{t}/(1+e^{t})$ is the well-known logistic function. With $f(\mathbf{x}_{i})=\beta_{0}+\mathbf{x}_{i}^{T}\boldsymbol{\beta}$, the negative log-likelihood for a single observation $(y_{i},\mathbf{x}_{i})$, often referred to as the logistic loss or log loss, is given by
$\mathcal{L}\left(f(\mathbf{x}_{i}),y_{i}\right)=\mathcal{L}(\beta_{0},\boldsymbol{\beta}\mid y_{i},\mathbf{x}_{i})=-y_{i}f(\mathbf{x}_{i})+\log(1+e^{f(\mathbf{x}_{i})}).$ (2)
This loss function arises directly from the Bernoulli likelihood where $\mathbb{P}(y_{i}=1|\mathbf{x}_{i})=p_{i}=S(f(\mathbf{x}_{i}))$ and $\mathbb{P}(y_{i}=0|\mathbf{x}_{i})=1-p_{i}=1-S(f(\mathbf{x}_{i}))=S(-f(\mathbf{x}_{i}))$. Minimizing the corresponding empirical loss (average negative log-likelihood)
$\mathcal{V}_{n}(f)=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}\left(f(\mathbf{x}_{i}),y_{i}\right)$ (3)
via standard maximum likelihood estimation requires $p<n$. When $p\geq n$, the estimation problem is typically ill-posed. Even when $p<n$, if $p$ is large relative to $n$, minimizing (3) often leads to overfitting of the training data and poor out-of-sample prediction accuracy.
@@
$\min_{\beta_{0}\in\mathbb{R},\,\boldsymbol{\beta}\in\mathbb{R}^{p}}\mathcal{V}_{n}(f)+\lambda P_{s}\left(\boldsymbol{\beta}\right),$ (4)
where $P_{s}(\boldsymbol{\beta})$ is a penalty function that induces sparsity in the coefficient vector $\boldsymbol{\beta}$. The tuning parameter $\lambda>0$ is usually determined in a data-driven way, typically by cross-validation (CV). A common and effective choice for the penalty term $P_{s}$ is the elastic net penalty (Zou and Hastie, 2005). It combines $\ell_{1}$ and $\ell_{2}$ regularization components:
$P_{s}(\boldsymbol{\beta})=\frac{1-\alpha}{2}\|\boldsymbol{\beta}\|_{2}^{2}+\alpha\|\boldsymbol{\beta}\|_{1},\quad\alpha\in[0,1].$ (5)
The mixing parameter $\alpha$ balances the sparsity-inducing $\ell_{1}$ norm $\|\boldsymbol{\beta}\|_{1}$ and the squared $\ell_{2}$ norm $\|\boldsymbol{\beta}\|_{2}^{2}$. This penalty includes Lasso (Tibshirani, 1996) ($\alpha=1$) and Ridge (Hoerl and Kennard, 1970) ($\alpha=0$) as special cases.
Next to Lasso and elastic net, some of the more popular sparse regularization methods are the adaptive Lasso (Zou, 2006), the relaxed Lasso (Meinshausen, 2007), the smoothly clipped absolute deviation (SCAD) estimator (Fan and Li, 2001) and the minimum concave penalized (MCP) estimator (Zhang, 2010). Additional approaches include Sure Independence Screening (SIS) (Fan and Lv, 2008), which first reduces dimensionality before applying a penalty like SCAD, and RuleFit (Friedman and Popescu, 2008), which combines sparse linear terms with rule-based terms derived from decision trees. A vast amount of asymptotic theory has been developed for a large class of regularized estimators, see e.g. Bühlmann and van de Geer (2011) for an extensive treatment. In summary, sparse regularization methods yield a single interpretable model with good prediction accuracy and extensive asymptotic theory that describes their behavior.
@@ 2.2 Ensemble Methods
The importance of diversity among the constituent models within an ensemble is a well-established principle, exemplified by popular methods like random forests (Breiman, 2001). In random forests, diversity is achieved by constructing individual decision trees using random subsets of candidate features at each split point. This strategy results in less correlated trees, which collectively enhance the ensemble’s predictive accuracy compared to using highly similar trees. Other techniques also leverage randomization to foster diversity. For instance, the random generalized linear models (RGLM) method of Song et al. (2013) combines bagging (Breiman, 1996a) (random sampling of data points) with the random predictor subspace method (Ho, 1998) (random sampling of features). This dual randomization creates varied training conditions for each base model, leading to a diverse set of intentionally simpler, decorrelated models. Sequential ensemble methods employ a different strategy. In methods like (extreme) gradient boosting (Chen and Guestrin, 2016), diverse trees are generated iteratively, with each new tree trained to correct the errors (residuals) made by the preceding ensemble. While individual trees might be relatively weak predictors on their own, this sequential error correction process builds a powerful and accurate final ensemble. A third approach to ensemble learning is stacking (Breiman, 1996b), where several different algorithms are trained on the same dataset, and their predictions are combined through a meta-learning algorithm that learns optimal weights for aggregating base model outputs. The specific relationship between various measures of diversity in classifier ensembles and their resulting prediction accuracy has been formally investigated by Kuncheva and Whitaker (2003).
In summary, many conventional ensemble methods face challenges regarding interpretability, as they often aggregate predictions from numerous relatively simple or “weak” base models, where the contribution of individual predictors can be obscured within the combined structure. While the overall ensemble might perform well, understanding the driving factors requires specific tools. Ad hoc measures have indeed been introduced to gauge predictor importance in some widely used ”black-box” ensemble methods, such as the Gini importance index commonly employed in random forests (Breiman, 2001). Finally, while ensemble methodologies often incorporate heuristic elements in their construction, theoretical understanding of their behavior has advanced. Significant research has been dedicated to analyzing their properties, leading to important results such as proofs of consistency for methods like Random Forests (Biau et al., 2008).
2.3 Split Logistic Regression Models